Jenny just went shoe shopping. Now she has 5 more pairs of shoes than her brother does. Together, they have 25 pairs of shoes.

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Jenny just went shoe shopping. Now she has 5 more pairs of shoes than her brother J + B = 25
J = B + 5
-------------
Substitute and solve for "B":
B+5 + B = 25
2B = 20
B = 10 (# of pairs her brother has)
J = B+5 = 15 (# Jenny has)
-----------
Cheers,
Stan H.
----------

Answer by ikleyn(52803)

.

It is a great sin to post a problem without a question . . .